Tickets for a concert were sold to adults for $3 and to students for $2. If the total receipts were $824 and twice as many adult tickets as student tickets were sold, then how many of each were sold?
Answer by bonster(299)

adult ticket sold=twice as many as student ticket sold
adult=2x
student=x

the adult ticket=$3 so (2x)(3)
student ticket=$2 so (x)(2)
(2x)(3)+2x=824
6x+2x=824
8x=824
x=103
if x=103, 2x=(2)(103)=206

103 student tickets and 206 adult tickets were sold